OpenStudy (anonymous):

An arithmetic sequence is represented by the explicit formula A(n) = 2 + 9(n - 1). What is the recursive formula?

OpenStudy (anonymous):

\[ a(2) - a(1)=9\\ a(3)-a(2)=9\\ a(4)-a(5)=9\\ \] Can you find the general pattern?

OpenStudy (anonymous):

a(n+1) = 9 + a(n)